# Policy Learning with Abstention

### v2

@@ Appendix C Missing proofs from Section˜4
Hence we can invoke Theorem 2.3 of kitagawa2018should, which establishes fast rates for policy learning under a (soft) margin condition. Our setting imposes a hard margin, which is a special case of theirs (formally, take the soft–margin parameter $\alpha\to\infty$). In fact, for a finite policy class, one may use the Bernstein inequality for each policy and union bound, instead of using uniform bounds for VC classes. We obtain using (16):
@@
$V(\pi^{B};{\mathcal{X}}_{rem})-V(\phi;{\mathcal{X}}_{rem})\lesssim\frac{D+\tfrac{1}{\delta}}{\kappa^{2}\,h\,n}$
@@
$r_{\tilde{\pi}}(x)$ $=\mathbf{1}\{\tilde{\pi}(x)\neq*\}[\tilde{\pi}(x)\mu_{1}(x)+(1-\tilde{\pi}(x))\mu_{0}(x)]$ (19)
$\quad+\mathbf{1}\{\tilde{\pi}(x)=*\}\cdot 0.5(\mu_{0}(x)+\mu_{1}(x));$ (20)
@@
$\min_{P_{X}(x)}\mathbb{E}_{P}[r_{\pi}(X)\mid X=x]$ $=\mathbf{1}\{\pi(x)\neq 0.5\}[\pi(x)\mu_{1}(x)+(1-\pi(x))\mu_{0}(x)-\alpha]$ (21)
$\quad+\mathbf{1}\{\pi(x)=0.5\}[0.5(\mu_{0}(x)+\mu_{1}(x))-\alpha/2]$ (22)
$=\mathbf{1}\{\tilde{\pi}(x)\neq 0.5\}[\tilde{\pi}(x)\mu_{1}(x)+(1-\tilde{\pi}(x))\mu_{0}(x)-\alpha]$ (23)
$\quad+\mathbf{1}\{\tilde{\pi}(x)=*\}[0.5(\mu_{0}(x)+\mu_{1}(x))-\alpha/2]$ (24)
@@
$\min_{P\in P_{\alpha}}V(\pi)$ $=\mathbb{E}[\mathbf{1}\{\tilde{\pi}(X)\neq*\}(\tilde{\pi}\mu_{1}+(1-\tilde{\pi})\mu_{0})+\mathbf{1}\{\tilde{\pi}(X)=*\}\cdot 0.5(\mu_{0}+\mu_{1})]$ (25)
$\quad-\alpha\cdot\mathbb{P}(\tilde{\pi}(X)\neq*)-(\alpha/2)\cdot\mathbb{P}(\tilde{\pi}(X)=*).$ (26)
@@ D.2 Safe Policy Learning
Baselines. We compare with several baseline algorithms for safe policy learning: (i) Safe EWM: direct EWM followed by comparing its estimated value with that of the baseline policy, and (ii) two versions (tt‑test LCB and clipped‑CI) of the HCPI algorithms proposed in thomas2015high.

### v3

@@ Appendix C Missing proofs from Section˜4
$\geq\mathbb{E}\Big[\mathbbm{1}\!\!\left\{{\widetilde{\pi}}(X)\neq*\right\}\left({\widetilde{\pi}}(X)Y(1)+(1-{\widetilde{\pi}}(X))Y(0)\right)\Big]+\mathbb{E}\left[\mathbbm{1}\!\!\left\{{\widetilde{\pi}}(X)=*\right\}\left(\frac{Y(1)+Y(0)}{2}+\frac{h}{2}\right)\right]$ ( margin condition (3))
@@
Hence we can invoke Theorem 2.3 of Kitagawa and Tetenov (2018), which establishes fast rates for policy learning under a (soft) margin condition. Our setting imposes a hard margin, which is a special case of theirs (formally, take the soft–margin parameter $\alpha\to\infty$). In fact, for a finite policy class, one may use the Bernstein inequality for each policy and union bound, instead of using uniform bounds for VC classes. We obtain using (13):
@@
$V(\pi^{B};{\mathcal{X}}_{rem})-V(\phi;{\mathcal{X}}_{rem})\lesssim\frac{D+\log\tfrac{1}{\delta}}{\kappa^{2}\,h\,n}$
@@
$r_{\tilde{\pi}}(x)$ $=\mathbf{1}\{\tilde{\pi}(x)\neq*\}[\tilde{\pi}(x)\mu_{1}(x)+(1-\tilde{\pi}(x))\mu_{0}(x)]$
$\quad+\mathbf{1}\{\tilde{\pi}(x)=*\}\cdot 0.5(\mu_{0}(x)+\mu_{1}(x));$
@@
$\min_{P_{X}(x)}\mathbb{E}_{P}[r_{\pi}(X)\mid X=x]$ $=\mathbf{1}\{\pi(x)\neq 0.5\}[\pi(x)\mu_{1}(x)+(1-\pi(x))\mu_{0}(x)-\alpha]$
$\quad+\mathbf{1}\{\pi(x)=0.5\}[0.5(\mu_{0}(x)+\mu_{1}(x))-\alpha/2]$
$=\mathbf{1}\{\tilde{\pi}(x)\neq 0.5\}[\tilde{\pi}(x)\mu_{1}(x)+(1-\tilde{\pi}(x))\mu_{0}(x)-\alpha]$
$\quad+\mathbf{1}\{\tilde{\pi}(x)=*\}[0.5(\mu_{0}(x)+\mu_{1}(x))-\alpha/2]$
@@
$\min_{P\in P_{\alpha}}V(\pi)$ $=\mathbb{E}[\mathbf{1}\{\tilde{\pi}(X)\neq*\}(\tilde{\pi}\mu_{1}+(1-\tilde{\pi})\mu_{0})+\mathbf{1}\{\tilde{\pi}(X)=*\}\cdot 0.5(\mu_{0}+\mu_{1})]$
$\quad-\alpha\cdot\mathbb{P}(\tilde{\pi}(X)\neq*)-(\alpha/2)\cdot\mathbb{P}(\tilde{\pi}(X)=*).$
@@ D.2 Safe Policy Learning
Baselines. We compare with several baseline algorithms for safe policy learning: (i) Safe EWM: direct EWM followed by comparing its estimated value with that of the baseline policy, and (ii) two versions (tt‑test LCB and clipped‑CI) of the HCPI algorithms proposed in Thomas et al. (2015).